OpenStudy (anonymous):

how do i find the missing side length using a^2+b^2=c^2 when the missing side length is b

OpenStudy (lgbasallote):

a^2 + b^2 = c^2 subtract a^2 from both sides b^2 = c^2 - a^2 take the square root of both sides \[b = \sqrt{c^2 - a^2}\]

OpenStudy (lgbasallote):

did you get the explanation @mariaha12 ?